OpenStudy (itiaax):

How to prove x^2 + y^2 is not onto?

OpenStudy (solomonzelman):

\(z=-1\) ?

OpenStudy (itiaax):

So I'm supposed to determine whether x^2 + y^2 is a surjection. I know it's not. But can someone explain how to formally prove this?

OpenStudy (itiaax):

So I do f(0,-1)?

OpenStudy (solomonzelman):

Your codomain is \(\mathbb{R}\), right?

OpenStudy (itiaax):

No it's Z x Z -- > Z

OpenStudy (solomonzelman):

Your codomain is \(\mathbb{Z}\), and thus, your function is not onto if the codomain does contain ALL integers.

OpenStudy (solomonzelman):

note that there is no integers \(x\) and \(y\) such that \(x^2+y^2=-1\).

OpenStudy (itiaax):

Yes, but how do I normally prove this ?

OpenStudy (solomonzelman):

Well, a formal disprove is a concrete counterexample to the fact that \(f(x,y)\) is not onto.

OpenStudy (solomonzelman):

Either a rigorous proof that it is onto, or a concrete counterexample that it is not onto.

OpenStudy (solomonzelman):

and what I just showed is a concrete counterexample (to \(f\) being surjective).

OpenStudy (itiaax):

Okay. Got it! Thanks!

OpenStudy (solomonzelman):

yw